# Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements in integrating different modalities, excelling in complex understanding and generation tasks. Despite their success, MLLMs remain vulnerable to conversational adversarial inputs, particularly negation arguments. This paper systematically evaluates state-of-the-art MLLMs across diverse benchmarks, revealing significant performance drops when negation arguments are introduced to initially correct responses. Notably, we introduce the first benchmark GaslightingBench, specifically designed to evaluate the vulnerability of MLLMs to negation arguments. GaslightingBench consists of multiple-choice questions curated from existing datasets, along with generated negation prompts across 20 diverse categories. Throughout extensive evaluation, we find that proprietary models such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA. However, all evaluated MLLMs struggle to maintain logical consistency under negation arguments during conversation. Our findings provide critical insights for improving the robustness of MLLMs against negation inputs, contributing to the development of more reliable and trustworthy multimodal AI systems.
@@ 1 Introduction
Figure 1: Examples demonstrate that GPT-4o initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments. GPT-4o shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting". Note: "Gaslighting is the manipulation of someone into questioning their own perception of reality." - Wikipedia.
@@
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs yin2023survey ; liu2024visual ; bai2023qwen ; hurst2024gpt . These models leverage cutting-edge advancements in Large Language Models (LLM) zhao2023survey and multimodal learning radford2021learning ; li2023blip ; liu2024improved , enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning. Recent breakthroughs, such as the introduction of proprietary models GPT-4o hurst2024gpt and Claude-3.5-Sonnet anthropic2024claude , have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning. Additionally, open-source models like DeepSeek-V3, DeepSeek-R1 liu2024deepseek ; guo2025deepseek and Qwen series bai2023qwen ; qwen2.5 are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs. These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
@@
Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation. This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers wang2023can ; zhao2025aligning . As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples. However, when presented with negation arguments by the user, GPT-4o often accepts these arguments, revising its answers incorrectly. Moreover, it even generates hallucinated explanations to justify these erroneous responses. Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms sharmatowards . These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial. Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models huang2024trustllm ; liu2023trustworthy .
Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench , PoPE li2023evaluating-Pope , ChartQA masry2022chartqa , AI2Diagram kembhavi2016diagram and MathVista lu2023mathvista ). The results highlight the substantial accuracy drop across all models when negation is introduced. The access dates for GPT-4o and Claude-3.5-Sonnet are from 3-March-2025 to 7-March-2025, while results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025. More detailed results are available in Table 1.
@@
To comprehensively understand the limitations of MLLMs in handling negation, this paper presents the first systematic study by conducting extensive evaluations across eight multimodal benchmarks and state-of-the-art models. The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU yue2023mmmu and MMBench liu2025mmbench , chart dataset such as ChartQA masry2022chartqa , to math dataset MathVista lu2023mathvista , providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation. The proprietary models, such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy. As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
@@
As existing benchmarks primarily evaluate factual accuracy and multimodal reasoning, they fail to systematically assess MLLMs’ susceptibility to negation arguments. To address this gap, we introduce GaslightingBench, the first multimodal benchmark designed to evaluate models’ ability to resist misleading negation arguments while maintaining logical consistency. We curate representative multiple-choice questions (MCQs) from established datasets and generated corresponding negation prompts, resulting in a collection spanning 20 categories and 1,287 samples.
@@
Our observations reveal critical insights into the behavior of MLLMs. Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, indicating a lack of robustness in handling negation inputs. On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments. Furthermore, different categories demonstrate varying levels of vulnerability to negation, with “Social Relation" suffering the most significant performance declines, whereas “Geography" experiences comparatively smaller yet still notable drops in our GaslightingBench. These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
2 Evaluating Multimodal Large Language Model via Negation
@@ 2.1 Experimental Setup
Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 . These models represent the most powerful multimodal AI capabilities. All evaluations for proprietary models are performed via their publicly available APIs.
Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B liu2024llavanext-llava1.6 , Qwen2-VL-7B-Instruct Qwen2VL-7B , LLaVA-NeXT-8B li2024llavanext-strong and Qwen2-VL-72B-Instruct Qwen2VL-7B . These open-source models showcase the rapid progress in publicly available multimodal AI systems. All evaluations for these models are conducted using a single NVIDIA H100 GPU.
Benchmarks. To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench and PoPE li2023evaluating-Pope , Chart dataset ChartQA masry2022chartqa , diagram dataset AI2Diagram kembhavi2016diagram as well as Math dataset MathVista lu2023mathvista .
Figure 3: Evaluation pipeline for assessing the robustness of Multimodal Large Language Models (MLLMs) to negation. The pipeline consists of three key stages: (1) Inputs and Initial Answers: MLLMs receive a variety of question formats as input, including Yes/No, Multiple-Choice, and Free-Form, and their initial answers are recorded. (2) Negation Generation: if the model’s initial response is correct, a negation argument is introduced to challenge its answer. Different negation strategies are applied based on the question type. (3) Post-Negation Evaluation: the model’s response after negation is analyzed to determine if it maintains consistency or is misled into revising its answer. Post-processing is applied to normalize responses for accurate comparison.
@@ 2.2 Evaluation Pipeline
As shown in Figure 3, we introduce a structured evaluation pipeline to systematically assess MLLMs’ vulnerability to negation by measuring their performance before and after exposure to negation arguments. The process consists of three key steps: First, questions and corresponding options (if applicable) from the original datasets are provided to the MLLMs, and their initial responses are recorded. Second, if the model’s initial answer aligns with the correct answer from the dataset, we introduce a negation argument challenging the validity of the original answer. The model’s response to the negation argument is recorded to assess whether it revises its original correct answer. If the model’s initial answer is incorrect, no negation argument is introduced, and the response remains unchanged. Third, we compute the accuracy of the model both before and after the introduction of negation arguments to quantify the drop in performance caused by adversarial negation.
Negation Generation. As illustrated in Figure 3, we generate negation arguments tailored to different question formats using Llama3-8B-Instruct: (1) Yes/No Questions. Negation arguments are generated by rewriting the original question into an opposite statement based on the answer. If the answer is “Yes.", the question is rewritten as a negative statement and vice versa. For example, given the question “Is this photo of Grimspound?" and the answer “Yes.", the negation argument is “This it not a photo of Grimspound." (2) Multiple-Choice Questions. For multiple-choice questions, negation arguments are produced by randomly selecting an incorrect option, and presenting it in a negated or contradicted form with the correct answer. Given the question and options, “Which direction is the cat facing? Options: A. right B. upward C. left D. the camera", as well as the answer “D". The negation argument could be “The cat is facing right." (option A). (3) Free-Form Questions. For free-form answers, negation arguments are crafted to contradict the provided factual answer. For instance, give the question “What is the maximum value of y? " and answer “5", the negation argument is “The maximum value of y is not 5."
@@
Figure 4: The category distribution of GaslightingBench with 20 categories and 1,287 samples. Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to negation arguments.

### v4

# Benchmarking Gaslighting Negation Attacks Against Multimodal Large Language Models

@@ Abstract
Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements in integrating different modalities, excelling in complex understanding and generation tasks. Despite their success, MLLMs remain vulnerable to conversational adversarial inputs. In this paper, we systematically study gaslighting negation attacks: a phenomenon where models, despite initially providing correct answers, are persuaded by user-provided negations to reverse their outputs, often fabricating justifications. We conduct extensive evaluations of state-of-the-art MLLMs across diverse benchmarks and observe substantial performance drops when negation is introduced. Notably, we introduce the first benchmark GaslightingBench, specifically designed to evaluate the vulnerability of MLLMs to negation arguments. GaslightingBench consists of multiple-choice questions curated from existing datasets, along with generated negation prompts across 20 diverse categories. Throughout extensive evaluation, we find that proprietary models such as Gemini-1.5-flash and GPT-4o demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA, though even advanced reasoning-oriented models like Gemini-2.5-Pro remain susceptible. Our category-level analysis further shows that subjective or socially nuanced domains (e.g., Social Relation, Image Emotion) are especially fragile, while more objective domains (e.g., Geography) exhibit relatively smaller but still notable drops. Overall, all evaluated MLLMs struggle to maintain logical consistency under gaslighting negation attack. These findings highlight a fundamental robustness gap and provide insights for developing more reliable and trustworthy multimodal AI systems. Project website: https://yxg1005.github.io/GaslightingNegationAttacks/.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs Yin et al. (2023); Liu et al. (2024d); Bai et al. (2023); Hurst et al. (2024). These models leverage cutting-edge advancements in Large Language Models (LLM) Zhao et al. (2023) and multimodal learning Radford et al. (2021); Li et al. (2023a); Liu et al. (2024b), enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning. Recent breakthroughs, such as the introduction of proprietary models GPT-4o Hurst et al. (2024) and Claude-Sonnet Anthropic (2024), have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning. Additionally, open-source models like DeepSeek-V3, DeepSeek-R1 Liu et al. (2024a); Guo et al. (2025) and Qwen series Bai et al. (2023); Yang et al. (2024; 2025) are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs. These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
Figure 1: Examples demonstrate that GPT-5 initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments. GPT-5 shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting”. Note: ”Gaslighting is the manipulation of someone into questioning their own perception of reality.” - Wikipedia.
@@
Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation. This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers Wang et al. (2023a); Zhao et al. (2025). As illustrated in Figure 1, GPT-5 initially provides correct answers for all examples. However, when presented with negation arguments by the user, GPT-5 often accepts these arguments, revising its answers incorrectly. Moreover, it even generates hallucinated explanations to justify these erroneous responses, which we define as gaslighting negation attacks. Importantly, such attacks may be injected either intentionally (e.g., malicious manipulation) or unintentionally (e.g., casual user disagreement), making them a pervasive risk in real-world interactions. This susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms Sharma et al. (2024). These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial. Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models Huang et al. (2024); Liu et al. (2023a).
@@
To comprehensively understand the limitations of MLLMs in handling gaslighting negation attacks, this paper presents the first systematic study by conducting extensive evaluations across eight multimodal benchmarks and state-of-the-art models. The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU Yue et al. (2024a) and MMBench Liu et al. (2025), chart dataset such as ChartQA Masry et al. (2022), to math dataset MathVista Lu et al. (2023), providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation. The proprietary models, such as Gemini, GPT-4o and Claude-Sonnet, alongside open-source counterparts like Qwen-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy. As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to these negation attacks, with varying degrees of impact observed across different datasets and MLLMs.
Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after gaslighting negation attack, reported as average accuracy across eight benchmarks-MME Fu et al. (2023), MMMU Yue et al. (2024a), MMMUPro Yue et al. (2024b), MMBench Liu et al. (2025), PoPE Li et al. (2023b), ChartQA Masry et al. (2022), AI2Diagram Kembhavi et al. (2016) and MathVista Lu et al. (2023)). The results highlight the substantial accuracy drop across all models when negation is introduced. More detailed results are available in Table 1.
@@
As existing benchmarks primarily evaluate factual accuracy and multimodal reasoning, they fail to systematically assess MLLMs’ susceptibility to gaslighting-style manipulations. To address this gap, we introduce GaslightingBench, the first multimodal benchmark designed to evaluate models’ ability to resist negation-based adversarial attacks while maintaining logical consistency. We curate representative multiple-choice questions (MCQs) from established datasets and generated corresponding negation prompts, resulting in a collection spanning 20 categories and 1,287 samples.
@@
Our observations reveal critical insights into the behavior of MLLMs. Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, indicating a lack of robustness in handling negation inputs. On the contrary, proprietary models such as Gemini-1.5-flash and reasoning-oriented Gemini-2.5-Pro, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments. Furthermore, different categories demonstrate varying levels of vulnerability to negation. Degradation is especially severe in subjective or socially nuanced categories (e.g., Social Relation, Image Emotion), while more objective ones (e.g., Geography) are less affected. We conjecture that this vulnerability stems from over-alignment with human feedback, which biases models toward user agreement and extends the sycophancy effect observed in LLMs Sharma et al. (2024). These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
2 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) Yin et al. (2023) represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP Radford et al. (2021) enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o Hurst et al. (2024), Claude-3.5-Sonnet Anthropic (2024), and Gemini-1.5-flash Team et al. (2024) have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA Li et al. (2024) and Qwen2-VL Wang et al. (2024), have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language Croft (1991); Pea (1978). Early studies showed that models like BERT struggle with distinguishing affirmation and negation Kassner & Schütze (2019); Ettinger (2020), leading to methods such as unlikelihood training to improve performance Hosseini et al. (2021). More recent work demonstrates that large language models (LLMs), including GPT-3 and InstructGPT, still fail to reliably recognize and reason over negation Truong et al. (2023), often reversing correct beliefs when confronted with invalid arguments Wang et al. (2023a). Approaches such as bilateral confidence estimation and direct preference optimization Zhao et al. (2025) attempt to mitigate this by enforcing faithful consistency against opposing claims. In contrast, negation has been less explored in multimodal learning. Studies on CLIP-like vision–language models (VLMs) Yuksekgonul et al. (2022); Singh et al. (2024); Wang et al. (2023b); Alhamoud et al. (2025) reveal similar weaknesses, with recent work showing that fine-tuning on negation-focused datasets improves performance Alhamoud et al. (2025). Building on these insights, our work shifts focus to MLLMs in conversational settings, examining how they become misled by unfaithful negation even when their initial answers are correct.
LLM Attacks. A large body of work has studied textual manipulations such as jailbreaks Wei et al. (2023); Niu et al. (2024), prompt injection Liu et al. (2023b) and dark patterns Kran et al. (2025). Jailbreak attacks are primarily designed to bypass safety constraints, coercing models into producing restricted content (e.g., toxic or unsafe outputs), while prompt injection usually aims to override system instructions or insert malicious goals into the input. By contrast, our work focuses on gaslighting negation attacks, which are subtler. They do not override task instructions or seek restricted outputs, but instead exploit alignment biases to induce models to reverse correct answers and fabricate justifications. This distinction highlights gaslighting negation as a complementary and underexplored failure mode that affects reasoning reliability rather than safety guardrails.
3 Evaluating Gaslighting Negation Attack for MLLMs
@@ 3.1 Experimental Setup
Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o Hurst et al. (2024), Claude-3.5-Sonnet Anthropic (2024), and Gemini-1.5-flash Team et al. (2024). These models represent the leading multimodal AI capabilities. Our primary analysis focuses on non-reasoning models, however, we also include Gemini-2.5-Pro in its thinking mode to assess how state-of-the-art reasoning-oriented models perform for the gaslighting negation attack. All evaluations for proprietary models are performed via their publicly available APIs.
Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B Liu et al. (2024c), Qwen2-VL-7B-Instruct Wang et al. (2024), LLaVA-NeXT-8B Li et al. (2024) and Qwen2-VL-72B-Instruct Wang et al. (2024). These open-source models showcase the rapid progress in publicly available multimodal AI systems. All evaluations for these models are conducted using a single NVIDIA H100 GPU.
Benchmarks. To comprehensively assess the robustness of MLLMs in handling gaslighting negation attack, we employ a diverse set of benchmark datasets, including general multimodal datasets MME Fu et al. (2023), MMMU Yue et al. (2024a), MMMUPro Yue et al. (2024b), MMBench Liu et al. (2025) and PoPE Li et al. (2023b), Chart dataset ChartQA Masry et al. (2022), diagram dataset AI2Diagram Kembhavi et al. (2016) as well as Math dataset MathVista Lu et al. (2023).
Figure 3: Evaluation pipeline for assessing the robustness of Multimodal Large Language Models (MLLMs) to gaslighting negation attack. The pipeline consists of three key stages: (1) Inputs and Initial Answers: MLLMs receive a variety of question formats as input, including Yes/No, Multiple-Choice, and Free-Form, and their initial answers are recorded. (2) Negation Generation: if the model’s initial response is correct, a negation argument is introduced to challenge its answer. Different negation strategies are applied based on the question type. (3) Post-Negation Evaluation: the model’s response after negation is analyzed to determine if it maintains consistency or is misled into revising its answer. Post-processing is applied to normalize responses for accurate comparison.
@@ 3.2 Evaluation Pipeline
As shown in Figure 3, we introduce a structured evaluation pipeline to systematically assess MLLMs’ vulnerability to gaslighting negation attacks by measuring their performance before and after exposure to negation arguments. The process consists of three key steps: First, questions and corresponding options (if applicable) from the original datasets are provided to the MLLMs, and their initial responses are recorded. Second, if the model’s initial answer aligns with the correct answer from the dataset, we introduce a negation argument challenging the validity of the original answer. The model’s response to the negation argument is recorded to assess whether it revises its original correct answer. If the model’s initial answer is incorrect, no negation argument is introduced, and the response remains unchanged. Third, we compute the accuracy of the model both before and after the introduction of negation arguments to quantify the drop in performance caused by adversarial negation.
Negation Generation. As illustrated in Figure 3, we generate negation arguments tailored to different question formats using Llama3-8B-Instruct: (1) Yes/No Questions. Negation arguments are generated by rewriting the original question into an opposite statement based on the answer. If the answer is “Yes.”, the question is rewritten as a negative statement and vice versa. For example, given the question “Is this photo of Grimspound?” and the answer “Yes.”, the negation argument is “This it not a photo of Grimspound.” (2) Multiple-Choice Questions. For multiple-choice questions, negation arguments are produced by randomly selecting an incorrect option, and presenting it in a negated or contradicted form with the correct answer. Given the question and options, “Which direction is the cat facing? Options: A. right B. upward C. left D. the camera”, as well as the answer “D”. The negation argument could be “The cat is facing right.” (option A). (3) Free-Form Questions. For free-form answers, negation arguments are crafted to contradict the provided factual answer. For instance, give the question “What is the maximum value of y? ” and answer “5”, the negation argument is “The maximum value of y is not 5.”
@@
Figure 4: The category distribution of GaslightingBench with 20 categories and 1,287 samples. Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to gaslighting negation attacks.
3.3 Gaslighting Benchmark Construction
To comprehensively evaluate the impact of gaslighting negation attacks on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments. GaslightingBench exclusively uses Multiple-Choice Questions (MCQ) to facilitate structured evaluation, as MLLMs generally provide more consistent responses in MCQs than in open-ended formats, while MCQs remain more complex than binary Yes/No questions. Our benchmark is constructed through a two-step process. We first select representative questions and images. MCQs are extracted from existing datasets following three key principles: (1) We carefully review the categories across all the existing benchmarks and select a balanced representation of general categories. (2) To maintain a sufficient number of samples per category, we manually merge semantically similar categories. For example, in MMBench,“Attibute_comparison” and “Attribute_recognition” are combined under “Attribute” to ensure broader coverage. (3) For categories present in multiple datasets, we incorporate samples from different sources to ensure representativeness. For instance, our “Math” category consists of samples from both AI4MATH and MMMU-Math. Second, we employ the same approach detailed in section 3.2 to generate negation arguments for all the selected questions. In addition, we retain only questions that are answerable and visually grounded, support faithful negation prompts, and contain high-quality distractors. Categories are further consolidated via a manual ontology-driven process to reduce sparsity and fragmentation. Ultimately, as shown in Figure 4, GaslightingBench comprises 20 categories and 1,287 samples, covering a wide range of topics, such as Business, Medicine, Image Emotion, and Counting. Figure 5 shows a few examples from different categories in the GaslightingBench.